OpenStudy (anonymous):

-4(2x - 1) = -6(x + 2) - 2

OpenStudy (fibonaccichick666):

So, what do you think? WHat have you tried so far?

OpenStudy (anonymous):

distribute -4 -6

OpenStudy (anonymous):

@Jamierox4ev3r please come

OpenStudy (anonymous):

ok

HanAkoSolo (jamierox4ev3r):



HanAkoSolo (jamierox4ev3r):

let me know if that makes sense @BabyPhoenix

OpenStudy (fibonaccichick666):

perfect, so what do you get when you distribute?

OpenStudy (anonymous):

8x+4=-6x-14

OpenStudy (anonymous):

14x+4=-14

OpenStudy (anonymous):

flutter all of u for not busting down on anwsers!!!!!!

OpenStudy (anonymous):

its -5

OpenStudy (anonymous):

*flutter

OpenStudy (anonymous):

14x=-18

OpenStudy (anonymous):

14x/14=-18/14 x=-18/14